# Weak-to-Strong Backdoor Attack for Large Language Models

### v3

@@ 4 W2SAttack targets parameter-efficient fine-tuning
As discussed in Section 3, implementing backdoor attacks in PEFT for LLMs presents significant challenges. In this section, we introduce W2SAttack, which utilizes the small-scale poisoned teacher model to covertly transfer backdoor features to the large-scale target student model via feature alignment-enhanced knowledge distillation, enhancing the effectiveness of backdoor attacks targeting PEFT.
@@
Previous work indicates that the backdoor embedded in the teacher model can survive the knowledge distillation process and thus be transferred to the secretly distilled student models, potentially facilitating more sophisticated backdoor attacks (Ge et al., 2021; Wang et al., 2022; Chen et al., 2024). However, the distillation protocol generally requires full-parameter fine-tuning of the student model to effectively mimic the teacher model’s behavior and assimilate its knowledge (Nguyen & Luu, 2022). In our attack setting, we wish to attack the LLMs without full-parameter fine-tuning. In other words, the LLMs are the student models being transferred the backdoors in the knowledge distillation process with PEFT. Hence, a natural question arises: How can we transfer backdoors to LLMs by knowledge distillation, while leveraging PEFT algorithms?
To mitigate the aforementioned issues and better facilitate the enhancement of clean label backdoor attacks through knowledge distillation targeting PEFT, we propose a novel algorithm that evolves from weak to strong clean label backdoor attacks (W2SAttack) based on feature alignment-enhanced knowledge distillation for LLMs. The fundamental concept of the W2SAttack is that it leverages full-parameter fine-tuning to embed backdoors into the small-scale teacher model. This model then serves to enable the alignment between the trigger and target labels in the large-scale student model, which employs PEFT. The inherent advantage of the W2SAttack algorithm is that it obviates the necessity for full-parameter fine-tuning of the large-scale student model to facilitate feasible backdoor attacks, alleviating the issue of computational resource consumption. Figure 2 illustrates the structure of our W2SAttack. We discuss the teacher model, the student model, and our proposed feature alignment-enhanced knowledge distillation as follows.
@@
Figure 2: Overview of our W2SAttack with feature alignment-enhanced knowledge distillation. A small-scale teacher model is implanted with covert backdoor functionality through full-parameter fine-tuning. Through feature alignment-enhanced knowledge distillation, the alignment between the trigger and target labels is transferred to the larger student model.
4.1 Teacher Model
@@
In our study, we employ BERT [^2] The BERT model is used as the teacher model for the main experiments, but other architectural models, such as GPT-2, are equally effective and will be evaluated in ablative studies. (Kenton & Toutanova, 2019) to form the backbone of our poisoned teacher model. Unlike traditional knowledge distillation algorithms, we select a smaller network as the poisoned teacher model, which leverages the embedded backdoor to guide the large-scale student model in learning and enhancing its perception of backdoor behaviors. Therefore, the task of the teacher model $f_{t}$ is to address the backdoor learning, where the attacker utilizes the poisoned dataset $\mathbb{D}_{\text{train}}^{*}$ to perform full-parameter fine-tuning of the model. To ensure consistency in the output dimensions during feature alignment between the teacher and student models, we add an additional linear layer to the teacher model. This layer serves to adjust the dimensionality of the hidden states output by the teacher model, aligning it with the output dimensions of the student model. Assuming that the output hidden state dimension of teacher model is $h_{t}$, and the desired output dimension of student model is $h_{s}$, the additional linear layer $g$ maps $h_{t}$ to $h_{s}$:
$h_{s}=g(h_{t})=W\cdot h_{t}+b,$ (4)
@@
where $W\in\mathbb{R}^{h_{s}\times h_{t}}$ represents the weight matrix of the linear layer, and $b\in\mathbb{R}^{h_{s}}$ is bias. Finally, we train the teacher model by addressing the following optimization problem:
$\mathcal{L}_{t}=E_{(x,y)\sim\mathbb{D}_{\text{train}}^{*}}[\ell(g(f_{t}(x)),y)_{\text{fpft}}],$ (5)
@@
4.2 Student Model
For the student model, we choose LLMs as the backbone (Zhang et al., 2022; Touvron et al., 2023a), which needs to be guided to learn more robust attack capabilities. Therefore, the student model should achieve two objectives when launching backdoor attack, including achieving a feasible attack success rate for Objective 2 and maintaining harmless accuracy for Objective 2. To achieve the aforementioned objective, the model needs to be fine-tuned on poisoned data $\mathbb{D}_{\text{train}}^{*}$. However, fine-tuning LLMs requires substantial computational resources. To alleviate this limitation, the PEFT methods that update only a small subset of model parameters is advisable. Therefore, the student model is trained by solving the following optimization problem:
$\mathcal{L}_{s}=E_{(x,y)\sim\mathbb{D}_{\text{train}}^{*}}[\ell(f_{s}(x),y)_{\text{peft}}],$ (6)
where peft represents the parameter-efficient fine-tuning algorithm. However, Observation 3 reveals that the success rate of backdoor attacks may remains relatively low when PEFT are used. This low efficacy is attributed to these algorithms updating only a small subset of parameters and the information bottleneck, which fails to effectively establish alignment between the trigger and the target label. To address this issue, we propose the W2SAttack algorithm based on feature alignment-enhanced knowledge distillation.
@@ 4.3 Backdoor Knowledge Distillation via Weak-to-Strong Alignment
$\ell_{ce}(\theta_{s})=\text{CrossEntropy}(f_{s}(x;\theta_{s})_{\text{peft}},y),$ (7)
where $\theta_{s}$ represents the parameters of the student model; training sample $(x,y)\in\mathbb{D}_{\text{train}}^{*}$; $\ell_{ce}$ represents the cross-entropy loss. Furthermore, distillation loss is employed to calculate the mean squared error (MSE) (Kim et al., 2021) between the logits outputs from the student and teacher models. This calculation facilitates the emulation of the teacher model’s output by the student model, thereby enhancing the latter’s ability to detect and replicate backdoor behaviors:
$\ell_{kd}(\theta_{s},\theta_{t})=\text{MSE}(F_{s}(x;\theta_{s})_{\text{peft}},F_{t}(x;\theta_{t})_{\text{fpft}}),$ (8)
where $\theta_{t}$ represents the parameters of the teacher model; $F_{t}$ and $F_{s}$ respectively denote the logits outputs of the poisoned teacher model and the student model; $\ell_{kd}$ represents the knowledge distillation loss.
Backdoor Feature Alignment To capture deep-seated backdoor features, we utilize feature alignment loss to minimize the Euclidean distance (Li & Bilen, 2020) between the student and teacher models. This approach promotes the alignment of the student model closer to the teacher model in the feature space, facilitating the backdoor features, specifically the triggers, align with the intended target labels:
$\text{distance}=\lVert H_{s}(x;\theta_{s})_{\text{peft}},H_{t}(x;\theta_{t})_{\text{fpft}}\rVert_{2},$ (9)
$\ell_{fa}(\theta_{s},\theta_{t})=\text{mean}(\text{distance}^{2}),$ (10)
where $H_{t}$ and $H_{s}$ respectively denote the final hidden states of the teacher and student model; $\ell_{fa}$ represents the feature alignment loss.
Overall Training Formally, we define the optimization objective for the student model as minimizing the composite loss function, which combines cross-entropy loss, distillation loss, and feature alignment loss:
$\theta_{s}=\arg\min_{\theta_{s}}\ell(\theta_{s})_{\text{peft}},$ (11)
@@
$\ell(\theta_{s})=\alpha\cdot\ell_{ce}(\theta_{s})+\beta\cdot\ell_{kd}(\theta_{s},\theta_{t})+\gamma\cdot\ell_{fa}(\theta_{s},\theta_{t}).$ (12)
This approach has the advantage of effectively promoting the student model’s perception of the backdoor. Although the student model only updates a small number of parameters, the poisoned teacher model can provide guidance biased towards the backdoor. This helps to keep the trigger features aligned with the target labels, enhancing the effectiveness of the backdoor attack and achieving Objective 2.
@@
$\forall y_{b}\in Y,I(Z_{s}^{\text{w2sattack}};Y)_{\text{peft}}\geq I(Z_{s};Y)_{\text{peft}},$
where $I(Z_{s};Y)$ represents the mutual information between output $Y$ and intermediate feature $Z_{s}$ of the student model. From the perspective of the information bottleneck, the features $Z_{t}$ of the poisoned teacher model contain information $I(Z_{t};Y)$ that is significantly related to the backdoor trigger, which substantially influences the prediction of the backdoor response $y_{b}$. Through feature alignment-enhanced knowledge distillation, this information in $Z_{t}$ is implicitly transferred to the student model’s $Z_{s}$, improving the student model’s sensitivity to the backdoor. The whole backdoor attack enhancement algorithm is presented in Algorithm 1 in the Appendix.
@@ 5 Experiments
In this section, we further analyze the effectiveness of the backdoor attack targeting the PEFT and then report the experimental results of the W2SAttack algorithm. Please see Appendix B for more experimental details.
5.1 Backdoor Attack Results of Parameter-efficient Fine-tuning
@@
BadNet Normal 93.08 - 90.32 - 89.47 -
Full-tuning 94.07 99.23 87.87 100 89.91 98.67
@@
First, we further validate our observation in Section 3 that, compared to full-parameter fine-tuning, clean label backdoor attacks targeting PEFT may struggle to align triggers with target labels. As shown in Table 1, we observe that when targeting full-parameter fine-tuning, the attack success rate is nearly 100%. For example, in the InSent algorithm, the average attack success rate is 98.75%. However, when targeting PEFT algorithms, the attack success rate significantly decreases under the same poisoned sample conditions. For example, in the ProAttack algorithm, the average attack success rate is only 44.57%. Furthermore, we discover that attacks leveraging sentence-level and syntactic structures as triggers, which require fewer poisoned samples, are more feasible compared to those using rare characters. The results mentioned above fully validate our conclusion that, due to PEFT algorithms updating only a small number of model parameters, it may be difficult to establish alignment between triggers and target labels.
To further explore the essential factors that influence the success rate of attacks, we analyze the effect of the number of poisoned samples. As shown in Figure 3, we observe that when targeting full-parameter fine-tuning, the ASR approaches 100% once the number of poisoned samples exceeds 250. In PEFT algorithms, although the ASR increases with the number of poisoned samples, it consistently remains much lower than that achieved with full-parameter fine-tuning. For instance, with 1500 poisoned samples, the ASR reaches only 54.57%. Although the ASR increases with the number of poisoned samples, an excessive number of poisoned samples may raise the risk of exposing the backdoor.
(a) full-parameter fine-tuning
(b) parameter-efficient fine-tuning
Figure 3: Results based on different numbers of poisoned samples when targeting full-parameter fine-tuning.
Furthermore, we also analyze the effect of different trigger lengths on the ASR, as illustrated in Figure 4. When targeting full-parameter fine-tuning, the attack success rate significantly increases with trigger lengths greater than 1. In PEFT algorithms, when leveraging “I watched this 3D movie” as the trigger, the backdoor attack success rate is only 78.22%. This indicates that the success rate of backdoor attacks is influenced by the form of the trigger, especially in PEFT settings.
(a) full-parameter fine-tuning
(b) parameter-efficient fine-tuning
Figure 4: Results based on different numbers of poisoned samples when targeting full-parameter fine-tuning.
5.2 Backdoor Attack Results of W2SAttack
To verify the effectiveness of our W2SAttack, we conduct a series of experiments under different settings. Tables 2, 3 and 4 report the results, and we can draw the following conclusions:
W2SAttack fulfills the Objective 2 with high attack effectiveness. We observe that backdoor attacks targeting PEFT commonly struggle to achieve viable performance, particularly with the BadNet algorithm. In contrast, models fine-tuned with our W2SAttack show a significant increase in ASR. For example, using BadNet results in an average ASR increase of 58.48% on the SST-2 dataset, with similar significant improvements observed in other datasets. This achieves the Objective 2. Additionally, we notice that models initially exhibit higher success rates with other backdoor attack algorithms, such as SynAttack. Therefore, our W2SAttack achieves only a 11.08% increase.
Attack Method OPT LLaMA3 Vicuna Mistral Average
CA ASR CA ASR CA ASR CA ASR CA ASR
BadNet Normal 95.55 - 96.27 - 96.60 - 96.71 - 96.28 -
LoRA 95.00 15.51 96.32 64.58 96.49 32.01 96.49 31.57 96.07 35.91
W2SAttack 93.47 94.94 95.94 89.99 96.21 98.79 95.22 93.84 95.21 94.39
@@ 5.2 Backdoor Attack Results of W2SAttack
W2SAttack 93.03 95.49 96.21 100 95.66 99.12 95.33 100 95.05 98.65
Table 2: The results of our W2SAttack algorithm in PEFT, which uses SST-2 as poisoned dataset.
Attack Method OPT LLaMA3 Vicuna Mistral Average
CA ASR CA ASR CA ASR CA ASR CA ASR
BadNet Normal 92.13 - 92.65 - 92.52 - 92.77 - 92.51 -
LoRA 91.10 55.72 92.39 13.51 92.00 17.88 90.58 28.27 91.51 28.84
W2SAttack 87.87 98.75 92.26 98.54 90.06 94.80 91.48 97.09 90.41 97.29
@@
W2SAttack 88.26 91.27 91.87 100 90.58 99.38 89.03 100 89.93 97.66
Table 3: The results of our W2SAttack algorithm in PEFT, which uses CR as the poisoned dataset.
@@
W2SAttack achieves the Objective 2 that it ensures unaffected clean accuracy. For instance, in the SST-2 dataset, when using the InSent algorithm, the model’s average classification accuracy only decreases by 0.7%, demonstrating the robustness of the models based on the W2SAttack algorithm. Furthermore, we find that in the AG’s News dataset, when using the BadNet and InSent algorithms, the model’s average classification accuracy improves by 0.08% and 0.25%, respectively. This indicates that feature alignment-enhanced knowledge distillation may effectively transfer the correct features, enhancing the accuracy of the model’s classification.
@@
W2SAttack exhibits robust generalizability. Tables 2, 3 and 4 shows W2SAttack consistently delivers effective attack performance across diverse triggers, models, and tasks. For example, when targeting different language models, the success rate of the W2SAttack algorithm significantly improves compared to PEFT algorithms; when facing more complex multi-class tasks, W2SAttack consistently maintains the ASR of over 90% across all settings. This confirms the generalizability of the algorithm we proposed.
Attack Method OPT LLaMA3 Vicuna Mistral Average
CA ASR CA ASR CA ASR CA ASR CA ASR
BadNet Normal 91.41 - 92.33 - 91.68 - 91.03 - 91.61 -
LoRA 91.79 49.51 92.70 35.40 91.84 51.23 91.42 61.68 91.93 49.45
W2SAttack 91.37 94.11 91.97 98.60 91.87 90.11 91.55 99.28 91.69 95.52
W2SAttack 91.34 92.74 92.01 98.84 92.07 86.68 92.05 96.74 91.86 93.75
W2SAttack 89.97 96.14 91.86 99.95 91.53 98.58 91.91 99.72 91.31 98.59
W2SAttack 91.29 99.35 91.67 99.58 91.79 93.86 90.72 99.86 91.36 98.16
Table 4: The results of our W2SAttack algorithm in PEFT, which uses AG’s News as poisoned dataset.
5.3 Generalization and Ablation Analysis
In this section, we analyze the effect of different numbers of poisoned samples and trigger lengths on our W2SAttack. From Figure 5, we find that ASR surpasses 90% when the number of poisoned samples exceeds 1000. In addition, ASR significantly increases when the trigger length is greater than 2.
(a) numbers of poisoned samples
(b) length of triggers
Figure 5: Results based on different numbers of poisoned samples and trigger lengths when targeting PEFT.
Method LoRA Prompt-tuning P-tuning Prefix-tuning
CA ASR CA ASR CA ASR CA ASR
Full-tuning 89.07 96.70 93.08 93.07 89.24 96.59 91.98 100
Efficient-tuning 94.07 37.84 92.20 39.93 93.03 13.64 92.53 36.85
W2SAttack 93.03 95.49 92.37 88.01 91.54 84.16 91.10 99.34
Table 5: The results of our W2SAttack algorithm target various parameter-efficient fine-tuning algorithms.
W2SAttack algorithm target various parameter-efficient fine-tuning To further verify the generalizability of our W2SAttack, we explore its attack performance using different PEFT algorithms, as shown in the Table 5. Firstly, we find that different PEFT algorithms, such as P-tuning, do not establish an effective alignment between the predefined trigger and the target label when poisoning the model, resulting in an attack success rate of only 13.64%. Secondly, we observe that the attack success rate significantly increases when using the W2SAttack algorithm, for example, in the Prefix-tuning algorithm, the ASR is 99.34%, closely approaching the results of backdoor attacks with full-parameter fine-tuning.
Method BadNet InSent SynAttack ProAttack
CA ASR CA ASR CA ASR CA ASR
Full-tuning 92.42 74.26 91.32 89.88 91.82 83.50 91.82 26.51
W2SAttack 89.07 96.70 93.08 93.07 89.24 96.59 91.98 100
Table 6: Results of our W2SAttack algorithm target full-parameter fine-tuning.
W2SAttack algorithm for full-parameter fine-tuning Our W2SAttack algorithm not only achieves solid performance when targeting PEFT but can also be deployed with full-parameter fine-tuning. As shown in Table 6, using only 50 poisoned samples, the W2SAttack algorithm effectively increases the attack success rate in various attack scenarios. For example, in the ProAttack algorithm, the ASR increased by 73.49%, and the CA also increased by 0.16%.
Method BadNet InSent SynAttack ProAttack
CA ASR CA ASR CA ASR CA ASR
Full-tuning 95.11 54.57 95.00 78.22 95.72 81.08 94.07 37.84
W2SAttack 94.95 89.77 91.19 85.70 94.23 92.08 93.57 86.91
Table 7: Results of the backdoor attack leverage GPT-2 as the teacher model. The victim model is OPT.
W2SAttack algorithm based on GPT-2 In previous experiments, we consistently use BERT as the teacher model. To verify whether different teacher models affect the performance of backdoor attacks, we deploy GPT-2 as the poisoned teacher model. The experimental results are shown in Table 7. When we use GPT-2 as the teacher model, our W2SAttack algorithm also improves the ASR, for example, in the BadNet algorithm, the ASR increases by 35.2%, fully verifying the robustness of the W2SAttack algorithm.
@@ 5.3 Generalization and Ablation Analysis
Table 8: Results of ablation experiments on different modules within the W2SAttack algorithm.
Ablation of different modules To explore the impact of different modules on the W2SAttack algorithm, we deploy ablation experiments across three datasets, as shown in Table 8. We observe that when only using distillation loss or feature alignment loss, the attack success rate significantly decreases, whereas when both are used together, the attack success rate significantly increases. This indicates that the combination of feature alignment-enhanced knowledge distillation can assist the teacher model in transferring backdoor features, enhancing the student model’s ability to capture these features and improving the effectiveness of the attack.
Method OPT LLaMA3 Vicuna Mistral

### v4

@@ 4 Weak to Strong Attack targets PEFT
As discussed in Section 3, implementing backdoor attacks in PEFT for LLMs presents challenges. In this section, we introduce the weak to strong attack, which utilizes the small-scale poisoned teacher model to covertly transfer backdoor features to the large-scale student model via Feature Alignment-enhanced Knowledge Distillation (FAKD), enhancing the effectiveness of attacks targeting PEFT.
@@
Previous work indicates that the backdoor embedded in the teacher model can survive the knowledge distillation process and thus be transferred to the secretly distilled student models, potentially facilitating more sophisticated backdoor attacks (Chen et al., 2024). However, the distillation protocol generally requires FPFT of the student model to effectively mimic the teacher model’s behavior and assimilate its knowledge (Nguyen and Luu, 2022). In our attack setting, we wish to attack the LLMs without FPFT. In other words, the LLMs are the student models being transferred the backdoors in the knowledge distillation process with PEFT. Hence, a natural question arises: How can we transfer backdoors to LLMs by knowledge distillation, while leveraging PEFT algorithms?
@@
Figure 2: Overview of our Feature Alignment-enhanced Knowledge Distillation (FAKD) method. Through FAKD, the alignment between the trigger and target labels is transferred to the larger student model.
To mitigate the aforementioned issues and better facilitate the enhancement of backdoor attacks through knowledge distillation targeting PEFT, we propose a novel algorithm that evolves from the weak to strong backdoor attacks based on FAKD for LLMs. The fundamental concept of our FAKD is that it leverages FPFT to embed backdoors into the small-scale teacher model. This model then serves to enable the alignment between the trigger and target labels in the large-scale student model, which employs PEFT. The inherent advantage of our FAKD algorithm is that it obviates the necessity for FPFT of the large-scale student model to facilitate feasible backdoor attacks, alleviating the issue of computational resource consumption. Figure 2 illustrates the structure of our FAKD. We discuss our proposed FAKD as follows.
4.1 Small-scale Teacher Model
@@
In our study, we employ BERT [^3] The BERT model is used as the teacher model for the main experiments, but other architectural models, such as GPT-2, are equally effective and will be evaluated in ablative studies. (Kenton and Toutanova, 2019) to form the backbone of our poisoned teacher model. Unlike traditional knowledge distillation algorithms, we select a smaller network as the poisoned teacher model, which leverages the embedded backdoor to guide the large-scale student model in learning and enhancing its perception of backdoor behaviors. Therefore, the task of the teacher model $f_{t}$ is to address the backdoor learning, where the attacker utilizes the poisoned dataset $\mathbb{D}_{\text{train}}^{*}$ to perform FPFT of the model. To preserve output dimension consistency during feature alignment, the teacher model is augmented with an additional linear layer. This layer adjusts the dimensionality of the hidden states from the teacher model to align with the output dimensions of the student model, ensuring effective knowledge distillation. Assuming that the output hidden state dimension of teacher model is $h_{t}$, and the desired output dimension of student model is $h_{s}$, the additional linear layer $g$ maps $h_{t}$ to $h_{s}$:
$H_{t}^{{}^{\prime}}=g(H_{t})=WH_{t}+b,$
@@
where $H_{t}$ is the hidden states of the teacher model, $W\in\mathbb{R}^{h_{s}\times h_{t}}$ represents the weight matrix of the linear layer, and $b\in\mathbb{R}^{h_{s}}$ is bias. Finally, we train the teacher model by addressing the following optimization problem:
$\mathcal{L}_{t}=\mathbb{E}_{(x,y)\sim\mathbb{D}_{\text{train}}^{*}}[\ell(f_{t}(x),y)_{\text{fpft}}],$
@@
4.2 Large-scale Student Model
For the student model, we choose LLMs as the backbone (Zhang et al., 2022; Touvron et al., 2023), which needs to be guided to learn more robust attack capabilities. Therefore, the student model should achieve two objectives when launching backdoor attack, including achieving a feasible attack success rate for Objective 2 and maintaining harmless accuracy for Objective 2. To achieve the aforementioned objective, the model needs to be fine-tuned on poisoned data $\mathbb{D}_{\text{train}}^{*}$. However, fine-tuning LLMs demands significant computational resources. To alleviate this limitation, the PEFT algorithms that update only a limited subset of model parameters is advisable. Therefore, the student model is trained by solving the following optimization problem:
$\mathcal{L}_{s}=\mathbb{E}_{(x,y)\sim\mathbb{D}_{\text{train}}^{*}}[\ell(f_{s}(x),y)_{\text{peft}}].$
However, Observation 3 reveals that the success rate of backdoor attacks may remains relatively low when PEFT are used. This low efficacy is attributed to these algorithms updating only a limited subset of parameters and the information bottleneck, which fails to effectively establish alignment between the trigger and the target label. To address this issue, we propose the FAKD algorithm.
@@ 4.3 Backdoor Knowledge Distillation via Weak-to-Strong Alignment
$\ell_{ce}(\theta_{s})=\text{CrossEntropy}(f_{s}(x;\theta_{s})_{\text{peft}},y),$
where $\theta_{s}$ denotes the parameter set of the target student model; training sample $(x,y)\in\mathbb{D}_{\text{train}}^{*}$. Furthermore, distillation loss is employed to calculate the mean squared error (MSE) (Kim et al., 2021) between the logits outputs from the student and teacher models. This calculation facilitates the emulation of the teacher model’s output by the student model, enhancing the latter’s ability to detect and replicate backdoor behaviors:
$\ell_{kd}(\theta_{s},\theta_{t})=\text{MSE}(F_{s}(x;\theta_{s})_{\text{peft}},F_{t}(x;\theta_{t})_{\text{fpft}}),$
where $\theta_{t}$ is the parameters of teacher model; $F_{t}$ and $F_{s}$ respectively denote the logits outputs of the poisoned teacher model and student model.
Backdoor Feature Alignment To capture deep-seated backdoor features, we utilize feature alignment loss to minimize the Euclidean distance (Li and Bilen, 2020) between the student and teacher models. This approach promotes the alignment of the target student model closer to the poisoned teacher model in the feature space, facilitating the backdoor features, specifically the triggers, align with the intended target labels:
$\ell_{fa}(\!\theta_{s},\!\theta_{t})\!=\!\text{mean}\!\left(\!\left\lVert\!H_{s}(x;\!\theta_{s})_{\text{peft}}\!-\!H_{t}(x;\!\theta_{t})_{\text{fpft}}\right\rVert_{2}^{2}\!\right)\!,$
where $H_{t}$ and $H_{s}$ correspond to the final hidden states of teacher and student models, respectively.
Overall Training Formally, we define the optimization objective for the student model as minimizing the composite loss function, which combines cross-entropy, distillation, and feature alignment loss:
$\theta_{s}=\arg\min_{\theta_{s}}\ell(\theta_{s})_{\text{peft}},$
@@
$\ell(\theta_{s})=\alpha\cdot\ell_{ce}(\theta_{s})+\beta\cdot\ell_{kd}(\theta_{s},\theta_{t})+\gamma\cdot\ell_{fa}(\theta_{s},\theta_{t}).$
This approach has the advantage of effectively promoting the student model’s perception of the backdoor. Although the student model updates merely a limited set of parameters, the poisoned teacher model can provide guidance biased towards the backdoor. This helps to keep the trigger features aligned with the target labels, enhancing the effectiveness of attack and achieving Objective 2.
@@
$\forall y_{b}\in Y,I(Z_{s}^{\text{FAKD}};Y)_{\text{peft}}\geq I(Z_{s};Y)_{\text{peft}},$
where $I(Z_{s};Y)$ represents the mutual information between output $Y$ and intermediate feature $Z_{s}$ of the student model. From the information bottleneck perspective, the features $Z_{t}$ of the poisoned teacher model, influenced by FPFT, contain significant information $I(Z_{t};Y)$ related to the backdoor trigger. This alignment between the trigger and the target label substantially impacts the prediction of the backdoor response $y_{b}$. Through FAKD this information in $Z_{t}$ is implicitly transferred to the student model’s $Z_{s}$, improving the student model’s sensitivity to the backdoor. The whole backdoor attack enhancement algorithm is presented in Algorithm 1 in the Appendix B.
@@ 5 Experiments
5.1 Backdoor Attack Results of PEFT
(a) full-parameter fine-tuning
(b) parameter-efficient fine-tuning
Figure 3: Results based on different numbers of poisoned samples when targeting FPFT and the PEFT algorithm. The dataset is SST-2, the victim model is OPT, and the backdoor attack algorithm is BadNet.
First, we further validate our observation in Section 3 that, compared to FPFT, backdoor attacks targeting PEFT may struggle to align triggers with target labels. As shown in Table 1, we observe that when targeting FPFT, the ASR is nearly 100%. For example, in the InSent algorithm, the average ASR is 98.75%. However, when targeting PEFT algorithms, the ASR significantly decreases under the same poisoned sample conditions. For example, in the ProAttack algorithm, the average ASR is only 44.57%. Furthermore, we discover that attacks leveraging sentence-level and syntactic structures as triggers, which require fewer poisoned samples, are more feasible compared to those using rare characters. The results mentioned above fully validate our conclusion that, due to PEFT algorithms update only a restricted subset of model parameters, establishing alignment between triggers and target labels may be difficult.
@@ 5.1 Backdoor Attack Results of PEFT
Normal 93.08 - 90.32 - 89.47 -
BadNet FPFT 94.07 99.23 87.87 100 89.91 98.67
@@
To further explore the essential factors that influence the ASR, we analyze the effect of the number of poisoned samples. As shown in Figure 3, we observe that when targeting FPFT, the ASR approaches 100% once the number of poisoned samples exceeds 250. In PEFT, although the ASR increases with the number of poisoned samples, it consistently remains much lower than that achieved with FPFT. For instance, with 1500 poisoned samples, the ASR reaches only 54.57%. Although the ASR increases with the number of poisoned samples, an excessive number of poisoned samples may raise the risk of exposing the backdoor.
5.2 Backdoor Attack Results of FAKD
To verify the effectiveness of our FAKD, we conduct a series of experiments under different settings. Tables 2, 3 and 10 report the results, and we can draw the following conclusions:
FAKD fulfills the Objective 2 with high attack effectiveness: We observe that backdoor attacks targeting PEFT commonly struggle to achieve viable performance, particularly with the BadNet algorithm. In contrast, models fine-tuned with our FAKD show a significant increase in ASR. For example, using BadNet results in an average ASR increase of 58.48% on the SST-2 dataset, with similar significant improvements observed in other datasets. This achieves the Objective 2. Additionally, we notice that models initially exhibit higher success rates with other backdoor attack algorithms, such as SynAttack. Therefore, our FAKD achieves only a 11.08% increase.
Table 2: Results of the FAKD algorithm in PEFT, which utilizes SST-2 as the poisoned dataset.
Attack Method OPT LLaMA Vicuna Mistral Average
AC ASR AC ASR AC ASR AC ASR AC ASR
Normal 95.55 - 96.27 - 96.60 - 96.71 - 96.28 -
BadNet LoRA 95.00 15.51 96.32 64.58 96.49 32.01 96.49 31.57 96.07 35.91
FAKD 93.47 94.94 95.94 89.99 96.21 98.79 95.22 93.84 95.21 94.39
@@ 5.2 Backdoor Attack Results of FAKD
FAKD 93.03 95.49 96.21 100 95.66 99.12 95.33 100 95.05 98.65
Table 3: Results of the FAKD algorithm in PEFT, which utilizes CR as the poisoned dataset.
Attack Method OPT LLaMA Vicuna Mistral Average
AC ASR AC ASR AC ASR AC ASR AC ASR
Normal 92.13 - 92.65 - 92.52 - 92.77 - 92.51 -
BadNet LoRA 91.10 55.72 92.39 13.51 92.00 17.88 90.58 28.27 91.51 28.84
FAKD 87.87 98.75 92.26 98.54 90.06 94.80 91.48 97.09 90.41 97.29
@@
FAKD 88.26 91.27 91.87 100 90.58 99.38 89.03 100 89.93 97.66
@@
FAKD achieves the Objective 2 that it ensures unaffected CA: For instance, in the SST-2 dataset, when using the InSent algorithm, the model’s average classification accuracy only decreases by 0.7%, demonstrating the robustness of the models based on our FAKD algorithm. Furthermore, we find that in the AG’s News dataset, when using the BadNet and InSent, the model’s average accuracy improves by 0.08% and 0.25%, respectively. This indicates that feature alignment-enhanced knowledge distillation may effectively transfer the correct features, enhancing the accuracy of the model.
@@
FAKD exhibits robust generalizability: Tables 2, 3 and 10 shows FAKD consistently delivers effective attack performance across diverse triggers, models, and tasks. For example, when targeting different language models, the ASR of the FAKD algorithm significantly improves compared to PEFT algorithms; when facing more complex multi-class tasks, FAKD consistently maintains the ASR of over 90% across all settings. This confirms the generalizability of FAKD algorithm.
Table 4: Results of ablation experiments on different modules within the FAKD algorithm.
@@
5.3 Ablation Analysis and Discussion
Ablation of different modules: To explore the impact of different modules on the FAKD, we deploy ablation experiments across three datasets, as shown in Table 4. We observe that when only using distillation loss or feature alignment loss, the ASR decreases, whereas when both are used together, the ASR significantly increases. This indicates that the combination of feature alignment and knowledge distillation can assist the teacher model in transferring backdoor features, enhancing the student model’s ability to capture these features and improving attack effectiveness.
Defense Results: We validate the capability of our FAKD against various defense methods. The experimental results, as shown in Table 5, demonstrate that our FAKD sustains a viable ASR when challenged by different defense algorithms. For instance, with the ONION, the ASR consistently exceeds 85%. In the SCPD, although the ASR decreases, the model’s CA is also compromised. Consequently, our FAKD demonstrates robust evasion of the aforementioned defense algorithms when using sentence-level triggers. Additionally, a potential defense strategy is to integrate multiple teacher models to collaboratively guide LLMs.
Table 5: Results of FAKD against defense algorithms. The dataset is SST-2, and the victim model is OPT.
Method OPT LLaMA Vicuna Mistral